OpenStudy (aaronandyson):

3250/x-5 - 3150/x = 5

OpenStudy (aaronandyson):

\[\frac{ 3250 }{ x-5 } - \frac{ 3150 }{ x } = 5\]

OpenStudy (aaronandyson):

@kohai

OpenStudy (anonymous):

subtract first

OpenStudy (anonymous):

\[\frac{3250x-3150(x-5)}{(x-5)x}=5\]

OpenStudy (anonymous):

then \[3250x-3150(x-5)=5x(x-5)\] divide both sides by \(5\) multiply out combine like terms solve the resulting quadratic

OpenStudy (aaronandyson):

x^2 - 25x = 3150 = 0 @satellite73

OpenStudy (phi):

\[ 3250x-3150(x-5)=5x(x-5) \\ 650x -630x+3150=x^2-5x \\ 20x+3150= x^2 -5x \\ x^2 -5x-20x -3150 = 0 \\ x^2 -25x -3150=0 \]

OpenStudy (phi):

factoring 3150 looks painful. try 70 as a factor.

OpenStudy (aaronandyson):

???